OpenStudy (anonymous):

Hello I have homework and I'm really confused 52) |x+3|=-2

OpenStudy (texaschic101):

| x + 3 | = -2 (x + 3) = -2 x = -2 - 3 x = -5 -(x + 3) = -2 -x - 3 = -2 -x = -2 + 3 -x = 1 x = -1 so x = -5 and x = -1 http://www.purplemath.com/modules/solveabs.htm

OpenStudy (texaschic101):

I left a link that I think you should look at.

OpenStudy (amistre64):

absolute value measures distance from an origin, so its never negative.

OpenStudy (amistre64):

at least it has no real solutions

OpenStudy (texaschic101):

check out the link...there are negative values for x

OpenStudy (amistre64):

we can graph |x+3| and -2, and try to see where they cross at: 

OpenStudy (amistre64):

can you test out your solutions?

OpenStudy (texaschic101):

| x + 3 | = -2 | -5 + 3 | = -2 | -2} = -2 2 = -2 oh....I see what you mean

OpenStudy (amistre64):

:) the confusion arises most likely from thinking that there has to be a solution ... when 'no real solutions' is a perfectly adequate way to define the solutions.

OpenStudy (texaschic101):

good explaining :)

OpenStudy (anonymous):

Can you guys halo me with another question 4|k+5|>8

OpenStudy (amistre64):

how far can you get on your own with it?

OpenStudy (amistre64):

divide and then split the absolute into 2 cases if need be.

OpenStudy (anonymous):

Sorry I don't understand what you just said.